Find the slope of the line through the given points.
1.(5,2),(7,10)
2.(-6,0),(0,-3)
3.(-5,-4),(-2,-9)

Answer by mananth(16949)

.(5,2),(7,10)
x1 y1 x2 y2
5 2 7 10

slope m = (y2-y1)/(x2-x1)
( 10 - 2 )/( 7 - 5 )
( 8 / 2 )
m= 4


2.(-6,0),(0,-3)
x1 y1 x2 y2
-6 0 0 -3

slope m = (y2-y1)/(x2-x1)
( -3 - 0 )/( 0 - -6 )
( -3 / 6 )
m= - 1/2
3.(-5,-4),(-2,-9)
Now you can walk through this problem